A wire of length 8x is bent into the shape of a square. Express the area A of the square as a function of x.
Answer by mananth(16946)

Perimeter of square = 8x
so length of side of square = 8x/4 = 2x
Area = L*W = 2x*2x = 4x^2 sq. units